OpenStudy (anonymous):

Find the inverse of the function. Enter NO INVERSE if no inverse exists. f (x) = 8x2 + 9

OpenStudy (anonymous):

x = 8y^2 + 9 (x-9)/8 = y^2 \[y = \pm \sqrt{\frac{x-9}{8}}\]

OpenStudy (anonymous):

so is that the answer or is there no inverse

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

yes answer or yes no inverse?

OpenStudy (anonymous):

the answer is NO INVERSE

OpenStudy (anonymous):

It could be because there are two functions

OpenStudy (anonymous):

ya we are dealing with one of one functions thank you